rxcklesskaisher:

The table below shows the values of y for different values of x: x y 0 0 1 3 2 6 3 9 Which equation shows the relationship between x and y? x = 3 + y y = 3x x = 3y y = 3 + x

jhonyy9:

 do you can continue for x = 4 , for x = 5 and for x = 6 ?

rxcklesskaisher:

its only up to x= 3 and y= 9

jhonyy9:

ok i know just to got more easy the right equation

rxcklesskaisher:

ok

jhonyy9:

when x = 1 how you got from 1 this y value of 3 ? or when x = 2 how you got from 2 the y value of 6 ?

jhonyy9:

how you can make from 2 the 6 ?

rxcklesskaisher:

times 3

jhonyy9:

exactly so x= 2 then for y=6 you need multiplie the 2 by 3 and for x= 1 how you get y= 3 ?

rxcklesskaisher:

2

jhonyy9:

what mean this 2 ?

rxcklesskaisher:

x2

jhonyy9:

how you make from 1 the y=3 ?

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @rxcklesskaisher x2 \(\color{#0cbb34}{\text{End of Quote}}\) 1×2 not equal 3

jhonyy9:

or when x=3 how you get the y=9 ?

rxcklesskaisher:

1x

rxcklesskaisher:

3

jhonyy9:

so 3 = 1×3 6 = 2×3 9 = 3×3 you know that given on this table x = 1 and 2 and 3 and y = 3 and 6 and 9 Which equation shows the relationship between x and y? x = 3 + y y = 3x x = 3y y = 3 + x

rxcklesskaisher:

y=3x

jhonyy9:

perfect

jhonyy9:

questions ?

rxcklesskaisher:

i choosed that one 2

jhonyy9:

ok - correct